# From Reliability to Security: How RIS-Assisted Adaptive SM and SSK Enhances Wireless Systems

### v1

@@ Abstract
This paper proposes two novel wireless transmission schemes, namely reconfigurable intelligent surface (RIS)-assisted received adaptive spatial modulation (RASM) scheme and RIS-assisted received adaptive space shift keying (RASSK) scheme, designed to enhance spectral efficiency (SE) and physical layer security (PLS). In both proposed schemes, transmitting bits are dynamically mapped at receive antennas by leveraging the characteristics of the RIS in each time slot, which enables the enhancement of signal-to-noise ratio (SNR) at specific selected antennas with near few power, thus leading a reliable and green wireless communication. This adaptive approach facilitates the conveyance of extra bits to the receiver, which means it needs less cost of radio-frequency chains at transmitter while improving SE. Besides, the proposed schemes offer an inherent PLS security advantage, as the eavesdropper is unable to completely detect signals reflected from the RIS. To comprehensively evaluate the performance of the proposed RASM and RASSK schemes, this paper presents a detailed analytical performance of their spectral efficiency, detection complexity, bit error rate, and secrecy rate, which are accompanied by insightful findings and conclusions. Simulation and analytical results demonstrate the superiority of the proposed schemes, showcasing their improved error performance and robustness against wiretapping, while also highlighting the potential of the RASM and RASSK schemes for future wireless applications.
@@ I-A Index Modulation
With the exponential growth of devices accessing wireless networks, communication throughput and data rates have surged recently. In response, novel techniques have emerged, with index modulation (IM) standing out. Numerous variants exist in the IM area. Prominent multi-antenna techniques include spatial modulation (SM), space shift keying (SSK), generalized space shift keying (GSSK), and others. These traditional schemes map antenna indices to convey information: single-antenna activation in SM/SSK and multi-antenna combinations in generalized schemes, enabling extra bits and enhancing spectral efficiency (SE) [1]. In IM schemes, constellation symbols modulated by $M$-ary phase shift keying (PSK) or quadrature amplitude modulation (QAM) are transmitted, while additional bits are encoded via antenna indices, each representing distinct transmitted bits. Unlike SM, SSK omits $M$-ary symbols, reducing digital modulation complexity at transmitter and receiver. As noted in [2], SSK exhibits slightly better bit-error-rate (BER) performance than SM with perfect channel state information (CSI) at the receiver. However, conventional SSK [2] demands numerous antennas for higher SE, raising costs and degrading BER. GSSK, introduced in [3], alleviates antenna costs by activating a fixed number, but yields inferior BER compared to SM and SSK. Thus, another groundbreaking IM approach is called addaptive spatial modulation (ASM), which provides flexibility through variable activated transmit antennas forming diverse antenna combinations (ACs), achieving superior SE over traditional SM, SSK, etc. [4]. By minimizing transmit antennas or ACs, ASM’s BER outperforms other IM schemes at equivalent SE. ASM incorporates constellation symbols, increasing complexity relative to adaptive space shift keying (ASSK), which excludes them. To lower the processing complexity at transmitter, this paper also explores ASSK, based on SSK principles where antennas emit pure power, simplifying modulation. As a trade-off, ASM and ASSK enhance SE beyond traditional IM schemes but introduce challenges, such as elevated radio frequency (RF) chain costs and physical layer security (PLS) concerns. To address these, a novel technique is applied to ASM and ASSK for system improvement.
@@ I-B Reconfigurable Intelligent Surface
As a promising technique garnering considerable attention in recent years, reconfigurable intelligent surfaces (RIS) holds great potential for application in future beyond 5th generation (B5G) and 6th generation (6G) wireless communication systems. The RIS comprises a chip, referred to as the intelligent unit, which is connected to the transmitter. This intelligent unit receives information about phase adjustments and subsequently controls the angles of the reflection elements (REs) accordingly [5]. The REs, as fabricated from the special materials that can reflect electromagnetic waves [6], are arranged in a two-dimensional (2D) planar array [7] and can be individually controlled by the intelligent unit. According to it, a deployment feature in the RIS is provided, that is, an extension in coverage of wireless communication through altering the communication delivery into the desired directions [8]. Unlike the traditional relays, the RF chains and amplifiers are not required in the RIS, which can better improve the power utilization in communication. Therefore, through the RIS, the phase shift of transmitted signals can be adjusted with near few powers and the signal-to-noise ratio (SNR) can be further enlarged at receiver, which is named the passive beamforming [9]. By possessing these advantages, the RIS can achieve the IM scheme at receive antennas by adjusting the phase to achieve the passive beamforming at one or more specific antennas, in which it can more effectively address the aforementioned issues of the ASM and ASSK schemes mentioned above and provides a viable consideration of system design for the proposed scheme, as illustrated as follows.
@@ I-C RIS-assisted IM schemes
Given extensive research on reconfigurable intelligent surfaces (RIS), IM is increasingly integrated into its applications. First, RIS acts as a passive relay to enhance IM communication quality. In [10]-[11], an RIS-assisted IM scheme is proposed where SSK operates at transmit antennas, improving BER performance via RIS assistance, as termed TSSK herein for convenience. RIS-assisted SM schemes are explored in various state-of-the-art works [12]-[16], with SM at transmit antennas and RIS boosting receiver SNR, sa called the TSM in this paper. To reduce antennas, RIS-assisted GSSK is introduced in [12], adhering to GSSK activation principles at transmit antennas, though with inferior BER compared to TSSK and TSM. Recent advancements include diverse RIS-assisted IM variants operable not only at transmitters but also at receivers via antenna selection, referred to as received IM herein. As a trade-off in RIS-assisted systems, received IM offers advantages like superior BER, reduced transmitter complexity, and potential gains in SE over related schemes [17]-[21]. In [17], RIS-assisted RSM and RSSK are first proposed with antenna selection at receive antennas, yielding good BER but omitting transmitter-RIS channel due to AP design, with room for SE improvement. Subsequently, RIS-assisted RGSM and RGSSK schemes emerge in [18]-[19], enhancing SE via novel low-complexity detectors, albeit with poorer BER than RSM and RSSK. All these schemes employ fixed selected or activated antennas for information mapping, limiting flexible SE enhancement or dynamic antenna adjustment. To overcome this, we propose two novel schemes with a flexible adaptive antenna selection algorithm, enabling dynamic resource configuration for improved SE adaptability and surpassing fixed-antenna constraints. The RIS-assisted transmitted ASM scheme is termed TASM, with ASM at the transmitter.
@@ I-D Contributions
The existing RIS-assisted IM schemes still face several critical challenges, e.g., contradicting with the low-cost nature of wireless communications to multiple RF chains, limit SE adaptability in dynamic channels by fixing antenna activation patterns, low PLS in wireless applications, etc. To address these issues, we propose two novel RIS-assisted received IM schemes, termed the the RIS-assisted received ASM (RASM) and RIS-assisted RASSK schemes, which dynamically map information bits to receive ACs instead of transmit antennas. Note that prior works on RIS-assisted IM schemes primarily focus on BER or data rate, overlooking the potential of RIS to enhance PLS in this domain. By leveraging RIS’s passive beamforming to enhance SNR at selected receive ACs, our schemes achieve higher SE, while the unique design in reflections inherently enhances the physical layer security against eavesdropping. Overall, the main contributions of this work can be briefly summarized as follows:
@@
I-E Notation
@@ II-A1 Applied Scenarios
This section illustrates the system models for the proposed RIS-assisted received IM schemes, RASM and RASSK, to enhance wireless networks. As depicted in Fig. 1, we consider a SIMO uplink in an IoT scenario where fixed sensors collect environmental data, e.g., temperature, and transmit it to a base station (BS). The sensor employs a single transmit antenna ($N_{t}=1$), while the BS has $N_{r}$ receive antennas. Due to obstructions, the direct link is blocked, relying solely on the RIS path. The RIS comprises $N$ REs and an intelligent unit linked to the transmitter. For instance, in the RASM scheme, bit sequence 010 maps to an AC selecting the first and fourth receive antennas; the intelligent unit directs the first RE subset to amplify the signal toward the first antenna and the second subset toward the fourth. Note that RASM transmits $M$-ary constellation symbols, whereas RASSK conveys only pure power, necessitating more receive antennas in RASSK. We denote $N_{r}^{1}$ for RASM and $N_{r}^{2}$ for RASSK, with $N_{r}^{1}\ll N_{r}^{2}$.
@@ II-A3 Specific Antenna Settings
For convenience, we define the vector of receive antenna to represent the various selected antennas in the $r$-th vector of selected AC as $\boldsymbol{v}_{r}=\left[0,1,\cdots,1\right]^{\mathsf{T}}$, where $\boldsymbol{v}_{r}\in\mathbb{C}^{1\times N_{r}}$, 1 represents the selected antennas, and 0 represents the non-selected ones. For example, based on mapping table in Fig. 1, if the transmitted bits sequence $r_{1}=001$ bits are transmitted, the selected antennas in the second AC are second third ones, of which the vector of receive antenna can be denoted as $\boldsymbol{v}_{2}=\left[0,1,1,0\right]^{\mathrm{T}}$. We can further define $v_{i,n}$ stands for the value of $i$-th element in $\boldsymbol{v}_{r}$ as well the $n$-th receive antenna at receiver. The total number of antennas in each AC can be represented as $N_{a}$, with $1\leq N_{a}\leq N_{r}$, which is not a fixed number in each time slot. For example, $N_{a}=1$ to represent bits sequence ’000’ in time slot 1, and $N_{a}=2$ to represent bits sequence ’001’ or ’010’ in time slot 2, as referring in mapping table in Fig. 1. This further enhances the flexiblity in improving SE and antenna configurations. The full set of ACs is given as $\varLambda(r)$, where $r=1,\mathrm{}2,\mathrm{}\cdots,J$ represents $r$-th selected AC and $J$ represents the total number of ACs. Thus, $J$ can be expressed as $J=\sum_{N_{a}}^{N_{r}}{\left(\begin{array}[]{c}N_{r}\\ N_{a}\\ \end{array}\right)}=2^{N_{r}}-1$. Based on the principle of ASM scheme, although $N_{r}$ receive antennas are equipped at receiver, only $N_{a}$ antennas can be selected to convey transmitted signals.
@@ II-A4 Working Principle
The REs are evenly divided into $N_{a}$ parts, which has $N_{E}=\lfloor\frac{N}{N_{a}}\rfloor$ number of REs, and each part is corresponding for one selected receive antenna to achieve the passive beamforming.  [^1] This even division strategy for RE allocation ensures fair resource distribution without real-time optimization. It enables straightforward passive beamforming by assigning equal RE subsets to boost SNR at each selected antenna. The method reduces computational overhead at the intelligent unit, suiting dynamic IoT scenarios with limited processing power. Besides, each antenna not only can receive the signals reflected from the REs that are corresponding for itself but also from the non-corresponding ones. Specifically, as shown in Fig.1, the first receive antenna can receive the signals reflected from the REs corresponding for itself (as shown in the red dashed line) and can be affected by those for the fourth receive antenna (as shown in blue dashed line) in the meanwhile, which leads to the constructive part and non-constructive part in the signals for $m$-th receive antenna, respectively. Notably, all $N_{r}$ receive antennas remain active during transmission, unlike traditional activation-based IM where unselected antennas are deactivated. Instead, the RIS employs passive beamforming to focus gain through constructive reflections on selected antennas while imposing non-constructive interference on others. The RE allocation strategy involves sequential assignment based on the sorted order of selected receive antenna indices in the AC, pre-agreed upon by transmitter and receiver. As shown in the mapping table in Fig. 1, for bit sequence 010 (mapping to antennas 1 and 4), the first RE part amplifies the signal toward antenna 1, and the second toward antenna 4. Similarly, for 100 (mapping to antennas 1, 2, and 3), the first part targets antenna 1, the second antenna 2, and the third antenna 3. This sequential approach enables straightforward implementation at the intelligent unit, which parses the bit sequence to identify the AC and sorts indices before phase adjustment. It ensures fairness by allocating equal RE resources to each selected antenna, avoiding complex optimization and maintaining low overhead in dynamic environments.

### v2

# arXiv:2512.03518v2[cs.IT] 02 Mar 2026

@@ Abstract
This paper proposes two novel wireless transmission schemes, namely reconfigurable intelligent surface (RIS)-assisted received adaptive spatial modulation (RASM) scheme and RIS-assisted received adaptive space shift keying (RASSK) scheme, designed to enhance spectral efficiency (SE) and physical layer security (PLS). In both proposed schemes, transmitting bits are dynamically mapped at receive antennas by leveraging the characteristics of the RIS in each time slot, which enables the enhancement of signal-to-noise ratio (SNR) at specific selected antennas with nearly few power, thus leading to a reliable and green wireless communication. Unlike conventional fixed-antenna RIS-RSM/GSSK, the term “adaptive” indicates the number of active antennas dynamically changes per symbol, conveying extra spatial information to break existing spectral efficiency bottlenecks. This adaptive approach facilitates the conveyance of extra bits to the receiver, which means it needs less cost of radio-frequency chains at transmitter while improving SE. Besides, the proposed schemes offer an inherent PLS security advantage, as the eavesdropper is unable to completely detect signals reflected from the RIS. To comprehensively evaluate the performance of the proposed RASM and RASSK schemes, this paper presents a detailed analytical performance of their spectral efficiency, detection complexity, bit error rate, and secrecy rate, which are accompanied by insightful findings and conclusions. Simulation and analytical results demonstrate the superiority of the proposed schemes, showcasing their improved error performance and robustness against wiretapping, while also highlighting the potential of the RASM and RASSK schemes for future wireless applications.
@@ I-A Index Modulation
With the exponential growth of devices accessing wireless networks, communication throughput and data rates have surged recently. In response, novel techniques have emerged, with index modulation (IM) standing out. Numerous variants exist in the IM area. Prominent multi-antenna techniques include spatial modulation (SM), space shift keying (SSK), generalized space shift keying (GSSK), and others. These traditional schemes map antenna indices to convey information: single-antenna activation in SM/SSK and multi-antenna combinations in generalized schemes, enabling extra bits and enhancing spectral efficiency (SE) [1]. In IM schemes, constellation symbols modulated by $M$-ary phase shift keying (PSK) or quadrature amplitude modulation (QAM) are transmitted, while additional bits are encoded via antenna indices, each representing distinct transmitted bits. Unlike SM, SSK omits $M$-ary symbols, reducing digital modulation complexity at transmitter and receiver. As noted in [2], SSK exhibits slightly better bit-error-rate (BER) performance than SM with perfect channel state information (CSI) at the receiver. However, conventional SSK [2] demands numerous antennas for higher SE, raising costs and degrading BER. GSSK, introduced in [3], alleviates antenna costs by activating a fixed number, but yields inferior BER compared to SM and SSK. Thus, another groundbreaking IM approach is called adaptive spatial modulation (ASM), which provides flexibility through variable activated transmit antennas forming diverse antenna combinations (ACs), achieving superior SE over traditional SM, SSK, etc. [4]. By minimizing transmit antennas or ACs, ASM’s BER outperforms other IM schemes at equivalent SE. ASM incorporates constellation symbols, increasing complexity relative to adaptive space shift keying (ASSK), which excludes them. To lower the processing complexity at transmitter, this paper also explores ASSK, based on SSK principles where antennas emit pure power, simplifying modulation. As a trade-off, ASM and ASSK enhance SE beyond traditional IM schemes but introduce challenges, such as elevated radio frequency (RF) chain costs and physical layer security (PLS) concerns. To address these, a novel technique is applied to ASM and ASSK for system improvement.
@@ I-B Reconfigurable Intelligent Surface
As a promising technique garnering considerable attention in recent years, reconfigurable intelligent surfaces (RIS) holds great potential for application in future beyond 5th generation (B5G) and 6th generation (6G) wireless communication systems. The RIS comprises a chip, referred to as the intelligent unit, which is connected to the transmitter. This intelligent unit receives information about phase adjustments and subsequently controls the angles of the reflection elements (REs) accordingly [5]. The REs, as fabricated from the special materials that can reflect electromagnetic waves [6], are arranged in a two-dimensional (2D) planar array [7] and can be individually controlled by the intelligent unit. According to it, a deployment feature in the RIS is provided, that is, an extension in coverage of wireless communication through altering the communication delivery into the desired directions [8]. Recently, the state-of-the-art applications of RIS have expanded profoundly, integrating advanced architectures, such as multi-layer refracting and self-powered absorptive surfaces, into complex space-air-ground integrated networks and satellite-terrestrial systems to fundamentally boost both coverage reliability and endogenous physical layer security [9, 10, 11, 12]. Unlike the traditional relays, the RF chains and amplifiers are not required in the RIS, which can better improve the power utilization in communication. Therefore, through the RIS, the phase shift of transmitted signals can be adjusted with near few powers and the signal-to-noise ratio (SNR) can be further enlarged at receiver, which is named the passive beamforming [13]. By possessing these advantages, the RIS can achieve the IM scheme at receive antennas by adjusting the phase to achieve the passive beamforming at one or more specific antennas, in which it can more effectively address the aforementioned issues of the ASM and ASSK schemes mentioned above and provides a viable consideration of system design for the proposed scheme, as illustrated as follows.
@@ I-C RIS-assisted IM schemes
Given extensive research on reconfigurable intelligent surfaces (RIS), IM is increasingly integrated into its applications. First, RIS acts as a passive relay to enhance IM communication quality. In [14]-[15], an RIS-assisted IM scheme is proposed where SSK operates at transmit antennas, improving BER performance via RIS assistance, as termed TSSK herein for convenience. RIS-assisted SM schemes are explored in various state-of-the-art works [16]-[20], with SM at transmit antennas and RIS boosting receiver SNR, so-called the TSM in this paper. To reduce antennas, RIS-assisted GSSK is introduced in [16], adhering to GSSK activation principles at transmit antennas, though with inferior BER compared to TSSK and TSM. Recent advancements include diverse RIS-assisted IM variants operable not only at transmitters but also at receivers via antenna selection, referred to as received IM herein. As a trade-off in RIS-assisted systems, received IM offers advantages like superior BER, reduced transmitter complexity, and potential gains in SE over related schemes [21]-[25]. In [21], RIS-assisted RSM and RSSK are first proposed with antenna selection at receive antennas, yielding good BER but omitting transmitter-RIS channel due to AP design, with room for SE improvement. Subsequently, RIS-assisted RGSM and RGSSK schemes emerge in [22]-[23], enhancing SE via novel low-complexity detectors, albeit with poorer BER than RSM and RSSK. All these schemes employ fixed selected or activated antennas for information mapping, limiting flexible SE enhancement or dynamic antenna adjustment. To overcome this, we propose two novel schemes with a flexible adaptive antenna selection algorithm, enabling dynamic resource configuration for improved SE adaptability and surpassing fixed-antenna constraints. While our proposed schemes rely on coherent detection, acquiring the requisite cascaded CSI typically requires dedicated pilot-based training protocols. To circumvent this estimation overhead, non-coherent transmission designs present a promising alternative [26, 27, 28], which demonstrate significant potential in reducing system complexity while enhancing secure communications. Unlike classic SM/SSK that requires multiple power-hungry transmit RF chains and relies on random fading, receive-side RIS-IM uses a single-antenna transmitter and a RIS to artificially reshape the channel. This paradigm minimizes IoT transmitter costs and enhances security. Advancing this, our schemes introduce an adaptive receive-antenna activation mechanism, fundamentally breaking the SE bottlenecks of conventional fixed-antenna RIS-IM. Besides, differing from traditional received IM schemes that rely on a fixed number of active antennas and physically deactivate the rest, our proposed schemes keep all receive antennas active and utilize RIS passive beamforming to dynamically vary the number of selected antennas per time slot. This adaptive mechanism fundamentally breaks the SE bottlenecks of fixed-antenna designs while intrinsically avoiding the computational overhead of complex real-time channel optimization. The RIS-assisted transmitted ASM scheme is termed TASM, with ASM at the transmitter.
@@ I-D Contributions
The existing RIS-assisted IM schemes still face several critical challenges, e.g., contradicting with the low-cost nature of wireless communications to multiple RF chains, limit SE adaptability in dynamic channels by fixing antenna activation patterns, low PLS in wireless applications, etc. To address these issues, we propose two novel RIS-assisted received IM schemes, termed the RIS-assisted received ASM (RASM) and RIS-assisted RASSK schemes, which dynamically map information bits to receive ACs instead of transmit antennas. Note that prior works on RIS-assisted IM schemes primarily focus on BER or data rate, overlooking the potential of RIS to enhance PLS in this domain. By leveraging RIS’s passive beamforming to enhance SNR at selected receive ACs, our schemes achieve higher SE, while the unique design in reflections inherently enhances the physical layer security against eavesdropping. Overall, the main contributions of this work can be briefly summarized as follows:
@@
I-E Organization and Notation
Organization: The rest of the paper is organized as follows: In Sec. II, the system model of the RASM and RASSK schemes with introduction of detection is given. Sec. III provides the illustration of random ACs selection. The analytical performance of SE, BER performance, and detection complexity is presented in Sec. IV. Sec. V provides the system model of the RASM and RASSK schemes with the eavesdropper and the analytical performance in PLS of the proposed schemes by deriving SR. The theoretical and simulation results of the RASM and RASSK schemes and other related schemes with different conditions are shown in Sec. VI. Eventually, the conclusion is given in Sec. VII.
@@ II-A1 Applied Scenarios
This section illustrates the system models for the proposed RIS-assisted received IM schemes, RASM and RASSK, to enhance wireless networks. As depicted in Fig. 1, we consider a SIMO uplink in an IoT scenario where fixed sensors collect environmental data  [^1] Since both transmit sensors and the receive BS are fixed in this scenario, mobility-induced channel fluctuations are beyond the scope of this paper., e.g., temperature, and transmit it to a base station (BS). The sensor employs a single transmit antenna ($N_{t}=1$), while the BS has $N_{r}$ receive antennas. Due to obstructions, the direct link is blocked, relying solely on the RIS path. The RIS comprises $N$ REs and an intelligent unit linked to the transmitter. For instance, in the RASM scheme, bit sequence 010 maps to an AC selecting the first and fourth receive antennas; the intelligent unit directs the first RE subset to amplify the signal toward the first antenna and the second subset toward the fourth. Note that RASM transmits $M$-ary constellation symbols, whereas RASSK conveys only pure power, necessitating more receive antennas in RASSK. We denote $N_{r}^{1}$ for RASM and $N_{r}^{2}$ for RASSK, with $N_{r}^{1}\ll N_{r}^{2}$.
@@ II-A3 Specific Antenna Settings
For convenience, we define the receive antenna activation vector for the $r$-th AC as $\boldsymbol{v}_{r}=[v_{r,1},v_{r,2},\dots,v_{r,N_{r}}]^{\mathsf{T}}\in\{0,1\}^{N_{r}\times 1}$, where $v_{r,n}=1$ indicates that the $n$-th receive antenna is selected, and $v_{r,n}=0$ otherwise. For example, based on the mapping table in Fig. 1, if the transmitted bit sequence is ‘001’, the second and third antennas are activated. The corresponding receive antenna vector is thus denoted as $\boldsymbol{v}_{2}=[0,1,1,0]^{\mathsf{T}}$. The number of active antennas in each AC, denoted as $N_{a}$ ($1\leq N_{a}\leq N_{r}$), dynamically changes within each symbol period. For instance, as shown in Fig. 1, $N_{a}=1$ is employed to transmit the bit sequence ‘000’, whereas $N_{a}=2$ is utilized for the sequences ‘001’ and ‘010’. This dynamic activation mechanism significantly enhances the flexibility of antenna configurations and substantially improves the SE. The full set of permissible ACs is defined as $\varLambda$, containing $J$ possible combinations. Consequently, the total number of ACs can be expressed as $J=\sum_{n=1}^{N_{r}}\binom{N_{r}}{n}=2^{N_{r}}-1$. Following the principle of the proposed adaptive scheme, although $N_{r}$ receive antennas are deployed, a dynamically varying subset of $N_{a}$ antennas is selected to convey the spatial information.
@@ II-A4 Working Principle
The REs are evenly divided into $N_{a}$ parts, which has $N_{E}=\lfloor\frac{N}{N_{a}}\rfloor$ number of REs, and each part is corresponding for one selected receive antenna to achieve the passive beamforming.  [^2] This even division strategy for RE allocation ensures fair resource distribution without real-time optimization. It enables straightforward passive beamforming by assigning equal RE subsets to boost SNR at each selected antenna. The method reduces computational overhead at the intelligent unit, suiting dynamic IoT scenarios with limited processing power. Besides, each antenna not only can receive the signals reflected from the REs that are corresponding for itself but also from the non-corresponding ones. Specifically, as shown in Fig.1, the first receive antenna can receive the signals reflected from the REs corresponding for itself (as shown in the red dashed line) and can be affected by those for the fourth receive antenna (as shown in blue dashed line) in the meanwhile, which leads to the constructive part and non-constructive part in the signals for $m$-th receive antenna, respectively. Notably, all $N_{r}$ receive antennas remain active during transmission, unlike traditional activation-based IM where unselected antennas are deactivated. Instead, the RIS employs passive beamforming to focus gain through constructive reflections on selected antennas while imposing non-constructive interference on others. The RE allocation strategy involves sequential assignment based on the sorted order of selected receive antenna indices in the AC, pre-agreed upon by transmitter and receiver. As shown in the mapping table in Fig. 1, for bit sequence 010 (mapping to antennas 1 and 4), the first RE part amplifies the signal toward antenna 1, and the second toward antenna 4. Similarly, for 100 (mapping to antennas 1, 2, and 3), the first part targets antenna 1, the second antenna 2, and the third antenna 3. This sequential approach enables straightforward implementation at intelligent unit, which parses the bit sequence to identify the AC and sorts indices before phase adjustment. It ensures fairness by allocating equal RE resources to each selected AC, strictly avoiding the overhead associated with real-time dynamic channel optimization algorithms, ensuring high efficiency.